How much energy is released when 546.0 grams of water at 5.000C is heated to 87.50C under constant pressure of 1.00 atm?


1
Expert's answer
2021-03-29T06:05:14-0400

q = m c "\Delta T"

q = heat released

c = specific heat of water = 4.186 j/g⁰C

"\Delta T" = T2 - T1 = 87.50⁰C - 5⁰C = 82.50⁰C

m = 546 g

q = 546×4.186×82.50

q = 188,558.37 joules energy is Released